OpenStudy (anonymous):

Someone please help me!!!!!! Solve the equation medal rewarded The square root of x = the square root of x-5, plus one

OpenStudy (anonymous):

isnt this already solved sometime back?

OpenStudy (anonymous):

I don't get it

OpenStudy (anonymous):

Can you help me pleade

OpenStudy (anonymous):

\[\sqrt{x} = \sqrt{x-5} + 1\]

OpenStudy (anonymous):

Squaring both sides, \[x = x-5 + 1 - 2\sqrt{x-5}\]

OpenStudy (anonymous):

ok till here?

OpenStudy (anonymous):

Where'd you get the 5 +1

OpenStudy (anonymous):

Oh wait nevermind

OpenStudy (anonymous):

Ok you can continue

OpenStudy (anonymous):

\[x-x + 5-1 = 2\sqrt{x-5}\]

OpenStudy (anonymous):

\[4 = 2\sqrt{x-5} => 2 = \sqrt{x-5}\]

OpenStudy (anonymous):

ok till here?

OpenStudy (anonymous):

What's the arrow for

OpenStudy (anonymous):

@natth88 that arrow means "Implies" I.e. the result.

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

squaring both sides again - \[4 = x-5 => x = 9.\]

OpenStudy (anonymous):

Thanks soo much